OpenStudy (anonymous):

I think its B or D but im not really sure.. can someone help Which of the following options is an equivalent function to f(x) = 4(3)^2x? f(x) = 36x f(x) = 4(9)x f(x) = 144x f(x) = 4(6x)

TheSmartOne (thesmartone):

Hint: \(\sf a(b)^2x = a(b^2)^x \neq (a\times b^2)^x\)

TheSmartOne (thesmartone):

@trap.lawd ...?

OpenStudy (beginnersmind):

It would be helpful if you gave f(x) as a drawing. It's not clear if x is in the exponent or not

OpenStudy (danjs):

\[\large f(x)=4(3)^{2*x} ~=~4(3^2)^x\]

TheSmartOne (thesmartone):

it is not A

TheSmartOne (thesmartone):

look at my hint, and notice that it has an equal sign and also a not equal sign :P

OpenStudy (danjs):

Memorize those 3 or 4 exponent rules that are in a table in your book prolly...